OpenStudy (anonymous):

"a sphere with the area A and volume V will follow the equation \[A = k * V^{p}\] find k and p

OpenStudy (anonymous):

I tried to rewrite it as 4 \pi r^{2} = k * (\frac{4}{3} \pi r^{3})^{p} But it didn't help that much. I should be able to solve it without log, but I don't see how.

OpenStudy (anonymous):

A=4*pi*r^2 V=4/3*pi*r^3  k=3/r and p=1